(5/u)-2 = (15/U)+4
Answer by nerdybill(7384)

(5/u)-2 = (15/U)+4
.
I'm assuming both u's are the same...
.
(5/u)-2 = (15/u)+4
Multiplying both sides by u, we get:
5-2u = 15+4u
Adding 2u to both sides:
5 = 15+6u
-10 = 6u
-10/6 = u
-5/3 = u